PNWNAS 1996, Registration Form

September 21, Vancouver Museum, Vancouver, British Columbia, Canada

Please send completed Registration Form with payment (cheques payable to Simon Fraser University) to: Manfred Trummer (PNWNAS), Dept of Mathematics, Simon Fraser University, Burnaby, B.C., Canada, V5A 1S6

Name:

Institution:

Address:

City, State/Province:

ZIP/Postal Code:

Telephone:  (  )
FAX:  (  )

Email:

Fees. Registration fee includes coffee/tea, refreshments, and lunch.

US Dollar Canadian Dollar
Participants with grants^{1}  $15  $ 20
regular fee  $10  $13
students/postdocs  $8  $ 10
Dinner : [number of tickets:  ]  $10  $13
Fees: $  , Dinner Tickets: $  , TOTAL: $  .

We will issue a receipt for the conference fee. □ Check here, if you need a receipt for your dinner tickets.

1If your conference registration fee is reimbursed by a grant or by your institution/company, we kindly ask you to support this event by paying the larger amount.